OpenStudy (anonymous):

What is the solution to the equation 14 x – 3 = 21 ? Answer: A) x ≈ –2.133 B) x ≈ –1.846 C) x ≈ 3.867 D) x ≈ 4.154

OpenStudy (anonymous):

c.

OpenStudy (anonymous):

Thank you:)))

OpenStudy (anonymous):

Solve for x: \[14x=21+3\] \[x = 24/14\] \[x\approx1.7142\] E?